# custom.bib

### v1

@@ Abstract
While Rotary Position Embeddings (RoPE) for natural language performs well and has become widely adopted, its adoption for other modalities has been slower. Here, we introduce Lie group Relative position Encodings (LieRE) that goes beyond RoPE in supporting higher dimensional inputs. We evaluate the performance of LieRE on 2D and 3D image classification tasks and observe that LieRE leads to marked improvements in performance (up to 6%), training efficiency (3.5x reduction), data efficiency (30%) compared to the baselines of RoFormer, DeiT III, RoPE-Mixed and Vision-Llama.
@@ 1 Introduction
While the attention mechanism has achieved widespread use, especially as part of the transformer architecture, attention is invariant to the order of its inputs and requires another mechanism to capture positional information of input tokens [vaswani2017attention]. This has spurred a line of work in the subarea of positional encodings—methods of encoding positional information for attention mechanisms.
In particular, Rotary Position Encoding (RoPE) has emerged as a powerful technique for encoding relative positional information in transformer-based models [RoFormer2024Su]. RoPE’s ability to capture relative position information has made it a popular choice for open-source foundation models such as LLaMA. In particular, RoPE implicitly captures relative positions. When the token in position $x$ attends to a token in position $y$, the effect of RoPE depends on $x-y$.
Despite the success of RoPE in sequence tasks, it is designed for one-dimensional sequence data. This has resulted in slow adoption for modalities with higher dimensional data, such as 2D images, and 3D medical images or data that includes a temporal dimension like videos or Electronic Health Records. Furthermore works that do attempt attempt to generalize RoPE to these modalities typically do so with a method tailored specifically to the modality at hand.
Our work aims to answer the question of whether there is a single relative position encoding scheme that well works across modalities of varying dimensions. If possible, this would enable the use of a common model backbone across a much larger variety of tasks.
$R(k_{1}x)$$R(k_{1}y)$$R(k_{2}x)$$R(k_{2}y)$
(a) 2D RoPE
$\exp(A\begin{bmatrix}x\\ y\end{bmatrix})$
(b) LieRE
Figure 1: Rotation matrices for a token generated from an image patch. $x,y$ denote the location of the patch from which the token was generated.
@@ 1.1 Contributions
We introduce Lie Relative Encodings (LieRE), a mechanism that allows the attention mechanism to learn how to utilize relative spatial information in its inputs. We show that LieRE is effective, without modification, across modalities and input dimensionality by evaluating it on both 2D and 3D classification tasks. Beyond improving classification accuracy, LieRE also reduces the amount of computing and data required during training. On the CIFAR100 task, this translates to 3.5x fewer training steps to achieve the same accuracy as the DeiT III baseline and outperforming the DeiT baseline trained on the full set with only a 70% subset of the data. Furthermore, LieRE is simple to implement and adapt to new modalities, requiring only a tokenizer that also outputs a position in $\mathbb{R}^{d}$ in addition to a standard embedding. In order to aid the reproducibility of our results we will post our code on github.
@@ 1.2.1 Position Encodings
The development of various position encodings has been ongoing since the original development of the attention mechanisms. There are three broad classes of positional encodings: 1) absolute 2) relative and 3) contextual.
Absolute encodings generally operate on a per token-level, modifying the embedding of a token to encode the location of the token in the text or media. Classic methods such as sinusoidal embeddings of learned embeddings achieve this by adding either a learned or carefully designed vector to the embedding of the token before it is passed through the transformer [vaswani2017attention, devlin2019bert, dosovitskiy2020image]. Absolute position measures position with respect to an absolute reference, such as the start of the text or the top left corner of an image.
Relative embeddings are focused on discarding the absolute coordinate system of absolute position encodings and encoding the relative positions of two tokens. This is particularly useful for images, where translation invariance can improve performance, but also widely used for text as well. The original relative position encodings modified the attention mechanism between two tokens with the difference in the positions of two tokens [shaw2018self, liu2021swin, liu2022swinv2]. However, this incurs quadratic computational complexity as an additional operation for every pair of tokens. Rotary Position Encodings (RoPE) avoid this penalty by utilizing the commutativity and orthogonality of 2D rotations. The keys and queries are rotated by a set of 2D rotations before computing the inner product, avoiding the quadratic blowup of earlier relative position schemes. The angle of these rotations depends on the position of the original token. The algebraic properties of attention and 2D rotation matrices ensure that the attention mechanism only depends on the relative positions of the tokens when they attend to each other [RoFormer2024Su]. RoPE is quite widely used in open source LLMs including the PaLM, Llama and Mixtral models [touvron2023llama, chowdhery2023palm, jiang2024mixtral]. However, RoPE can perform poorly on inference for larger context sizes than the model was trained on.
@@ 1.2.2 Extensions of RoPE
One notable one is context extension, which aims to address the fact that models trained on short documents tend to perform poorly on long documents. Methods like NTK-aware context extension, YaRN and LongROPE focus on enabling already trained models to handle long context, both with and without finetuning [ding2024longrope, peng2023yarn].
Another line of work has been specifically focused on adapting RoPE to image tasks. Both VisionLlama and RoPE-Mixed present relative position encodings inspired by RoPE that are able to encode 2D positional data [chu2024visionllama, heo2024rotary] I key difference is that RoPE-Mixed has a learnable component, whereas VisionLlama does not.
@@ 2.1 Lie Groups in the Context of Attention
In this section, we aim to provide a minimal introduction to Lie groups so that the reader is able to understand the mathematical motivations behind LieRE. Lie groups are very well studied, especially in the context of representation theory and standard texts including [fulton2013representation] are able to provide a more extensive introduction to the subject.
In this context, Lie groups are smooth manifolds that are closed under matrix multiplication and inversion. For every Lie group, the matrix exponential provides a smooth bijective map from a subset of $\mathbb{R}^{n\times n}$ (hereto referred to as the generator set) to the Lie group. The exponential map is a diffeomorphism and has the following key property for generators $x,y$ close together:
$\exp(x-y)=\exp(-y+x)\approx\left(\exp(y)\right)^{-1}\exp(x)$ (1)
@@
$exp(x-y)=exp(x)exp(y)^{-1}=exp(y)^{-1}exp(x).$ (2)
Our work examines the tradeoff between using the stronger property in equation (2) or increased capacity and the weaker property (1). In short, we find that benefit of representational flexibility exceeds the benefit of exactly honoring equation (2).
@@ 3.1 Rotations for Relative Positions in Attention
We start by using a linear map to embed the positions as skew-symmetric matrices $P_{i}$ and $P_{j}$ and compute high dimensional rotations as $R_{i}=\exp(P_{i})$ and $R_{j}=\exp(P_{j})$.
Recall that, for every pair of tokens, the attention mechanism computes the inner products between their key and query vectors, $k_{i}^{T}q_{j}$. We encode the positions by multiplying by the rotation matrices in the prior part. In particular, we update the keys and queries as $k_{i}^{\prime}=k_{i}R_{i}$ and $q_{j}=q_{j}R_{j}$. This results in an updated inner product of $(R_{i}k_{i})^{T}R_{j}q_{j}=k_{i}R_{i}^{T}R_{j}q_{j}=k_{i}R_{i}^{-1}R_{j}q_{j}.$
Recall that by equation 1, $R_{i}^{-1}R_{j}=\exp(P_{i})^{-1}\exp(P_{j})\approx\exp(P_{j}-P_{i})$. In other words, the inner product automatically computes the relative position encoding. Note that this is exactly the mechanism behind the original RoPE paper. The difference is that LieRE uses the full set of high dimensional rotations and relaxes the constraint that equation 1 holds with equality. We make the description of the method more precise in the next section.
@@ 3.2 LieRE
LieRE modifies the attention mechanism by applying a rotation to the keys and queries before computing their inner products. However, where RoPE applies a block-diagonal $n\times n$ with 2D rotation matrices as blocks, LieRE applies a $n\times n$ rotation matrix. Using the fact that every rotation matrix can be represented as the matrix exponential of a skew-symmetric matrix, we parametrize the rotations with generators before the matrix exponential. This skew-symmetric generator is computed via a learned linear map from the positions to the generator space. Let $x\in\mathbb{R}^{n}$ indicate the position of the token, and $A:\mathbb{R}^{d}\rightarrow\text{Skew}_{n}(\mathbb{R})$ be a learnable linear map to the set of skew-symmetric matrices. At the start of every forward pass, we compute the LieRE rotation matrix for each token as
@@
Then, for each attention mechanism, the keys and queries for a particular token are updated as $\mathrm{key^{\prime}}:=R_{LieRE}\cdot\mathrm{key},\mathrm{query^{\prime}}:=R_{LieRE}\cdot\mathrm{query}$ We present attention with RoPE and LieRE side by side in algorithms 1 and 2.
Algorithm 1 RoPE Attention (prior work)
@@
5:         $X_{\text{rotated}}[i]\leftarrow X[i]\cdot\text{cos\_pos}_{i}-X[i+1]\cdot\text{sin\_pos}_{i}$
6:         $X_{\text{rotated}}[i+1]\leftarrow X[i]\cdot\text{sin\_pos}_{i}+X[i+1]\cdot\text{cos\_pos}_{i}$
7:     end for
8:     return $X_{\text{rotated}}$
9:end procedure
10:procedure RoPEAttention($Q,K,V$)
11:     $p\leftarrow\text{tokenPositions}$
[…]
13:     $K_{\text{rotated}}\leftarrow\textsc{RoPE}(K,p,d)$
14:     $Q_{\text{rotated}}\leftarrow\textsc{RoPE}(Q,p,d)$
15:     $\text{Attention}\leftarrow\text{softmax}(\frac{Q_{\text{rotated}}K_{\text{rotated}}^{T}}{\sqrt{d}})V$
16:     return Attention
17:end procedure
Algorithm 2 LieRE Attention (our work)
@@
8:     $K_{\text{rotated}}\leftarrow\textsc{LieRE}(K,p,A)$
9:     $Q_{\text{rotated}}\leftarrow\textsc{LieRE}(Q,p,A)$
10:     $\text{Attention}\leftarrow\text{softmax}\left(\frac{Q_{\text{rotated}}K_{\text{rotated}}^{T}}{\sqrt{\text{dim}(K)}}\right)V$
@@
Note that if we are applying our method to images it is possible to impose a sparsity structure on the generator that recovers RoPE-mixed without the absolute position embeddings [heo2024rotary].
@@ 4.1 Model and Training Parameters
In order to isolate the benefit of LieRE, we use the standard transformer backbone modified to be able to switch between relative position encoding types. The backbone is configured as ViT-B, with 12 layers, a hidden dimension of 768, and an intermediate dimension of 3096. All experiments use dropout of 0.1. All experiments use RandAugment [cubuk2020randaugment].
We avoid using pre-trained weights in order to maximize the comparability of results between methods. All trainers use ADAM with the default parameters for pytorch.
In order to ensure a fair comparison, we explicitly avoid tuning hyperparameters with LieRE and use the same default hyperparameters for all experiments (Appendix A).
@@ 4.2 Datasets and Tasks
Our experiments are designed to evaluate the efficacy of LieRE as a position encoding across a diverse set of modalities. We evaluate LieRE on the classification of 2D (images) and 3D (videos and medical images) data. For 3D data and ImageNet (2D), we focus on accuracy. For CIFAR100 (2D), where training is less resource intensive, we also evaluate LieRE’s data and training compute efficiency.
@@
For 2D data we evaluate performance on the CIFAR100 and ImageNet image classification task [krizhevsky2009learning, deng2009imagenet]. All hyperparameters and other details necessary for the reproducibility of experiments can be found in appendix A.
We partition our evaluation of performance into four parts. In the first part, we examine accuracy across a range of model architectures on both CIFAR100 and ImageNet. To ensure a fair comparison, we implement all positional encodings on the same [standard] transformer backbone. These experiments are meant to identify the best-performing method of encoding positions.
In the second part, we take advantage of the relatively modest amount of compute resources necessary to train a model for CIFAR 100 to examine LieREs impact on data efficiency. Here we limit our focus to LieRE, RoPE-Mixed, VisionLlama embeddings and absolute position embeddings. We train every model on 30%, 40%,…,90% subsets of CIFAR100 in order to evaluate the sensitivity of each model to the amount of training data. Since training starts to overfit quite early for the smaller datasets, we use the maximum validation accuracy for each run when comparing models.
We also measure training compute efficiency by comparing the number of training steps necessary to achieve a fixed level of validation accuracy.
In the third part, we evaluate the scaling laws of the LieRE generator capacity by imposing a block diagonal structure on the generator that allows us to vary the added capacity. With a head dimension of $64$ LieRE adds $4032$ parameters to a transformer designed for $2D$ data. While this is a small number of parameters, they lie in a critical point in the architecture, and understanding the marginal impact of this capacity can help inform computational tradeoffs.
The block diagonal generator structure results in a block-diagonal rotation matrix after the matrix exponential. The parameters for each block are learned separately. This setup allows us to vary the capacity of LieRE without modifying the rest of the model.
In the last part, we measure how much different models depend on the positional information in the image by shuffling the patches. A higher accuracy drop means the model relies more on the positions of the patches during inference.
@@
In this section, we introduce Rotary Position Encodings for 3D data and compare LieRE-based transformers with transformers with RoPE-mixed and absolute encoding similar to the previous section [arnab2021vivit, heo2024rotary].
For the 3D experiments we examine video classification performance in the UCF101 dataset [soomro2012ucf101] and hemorrhage classification on the RSNA dataset [rsna2019data]. We treat the RSNA dataset as a binary classification problem, asking the model to identify whether a 3D brain scan has a hemorrhage or not.
Again, we did not optimize any hyperparameters for the LieRE model. We additionally use Mixup described standard used augmentation method for videos [zhang2017mixup]. The full set of hyperparameters may be found in appendix A.

### v2

@@ Abstract
While Rotary Position Embeddings (RoPE) for large language models have become widely adopted, their application for other modalities has been slower. Here, we introduce Lie group Relative position Encodings (LieRE) that goes beyond RoPE in supporting n-dimensional inputs. We evaluate the performance of LieRE on 2D and 3D image classification tasks and observe that LieRE leads to marked relative improvements in performance (up to 9.7% for 2D and up to 25.5% for 3D), training efficiency (3.5x reduction), data efficiency (30%) compared to the baselines of DeiT III, RoPE-Mixed and Vision-Llama. https://github.com/Stanford-AIMI/LieRE
@@ 1 Introduction
While the attention mechanism has achieved widespread use, especially as part of the transformer architecture, attention is invariant to the order of its inputs and requires another mechanism to capture positional information of input tokens [vaswani2017attention]. This has spurred a line of work in the subarea of positional encodings—methods of encoding positional information in attention mechanisms.
In particular, Rotary Position Encoding (RoPE) has emerged as a technique for encoding relative positional information of text tokens in transformer-based models [RoFormer2024Su]. RoPE’s ability to capture relative position information has made it a popular choice for open-source language foundation models such as LLaMA. In particular, RoPE implicitly captures relative positions. When the token in position $p_{1}$ attends to a token in position $p_{2}$, the effect of RoPE depends on $p_{1}-p_{2}$.
Despite the success of RoPE in sequence tasks [touvron2023llama, chowdhery2023palm], it is designed for one-dimensional sequence data. This has resulted in slow adoption for modalities with higher dimensional data, such or data that includes a temporal dimension like videos.
Our work aims to answer whether a single position encoding scheme can work well across both 2D and 3D modalities. If possible, this would enable the use of a simpler common model backbone across a variety of tasks.
[…]
Figure 1: LieRE sketch, where $A$ is a learnable skew symmetric matrix and $R_{i}=\exp(A[x\;y\;z])\in\mathbb{R}^{d\times d}$ is the rotation matrix for the first patch in the top left corner of the input with the position $P_{i}=(x,y,z)\in\mathbb{R}^{N}$. $d$ is the head dimension. The upper triangle contains learnable $\textrm{LieRE}_{\theta}$ parameter, while the lower triangle contains their negatives, -$\textrm{LieRE}_{\theta}$, reflecting the skew-symmetric nature of the matrix.
[…]
Figure 2: Comparisons of basis matrix $A$ of our work, Rope-Mixed [heo2024rotary], VisionLlama [chu2024visionllama]. First row: learnable parameter (yellow), not learnable parameter (blue) in basis matrix. Second row: rotation matrices $\exp(Ax)$ shared across the stem (gray) or different for each layer and head (colorful). MHA := Multi-Head-Attention, FF := Feed Forward, L[index] := Layer index, RxH := Rotation Matrix per Head, $\textrm{LieRE}_{\theta}$ := the number of LieRE parameters.
@@ 1.1 Contributions
We introduce Lie Relative Encodings (LieRE), a mechanism that allows the attention mechanism to learn how to utilize relative spatial information of its inputs. We show that LieRE is effective on both 2D and 3D inputs of various modalities. Beyond improving classification accuracy, LieRE also reduces the amount of compute and data required during training to achieve a fixed accuracy. On the CIFAR100 task, this translates to 3.5x fewer training steps to achieve the same accuracy as the DeiT baseline and outperforming the DeiT baseline trained on the full set with only a 70% subset of the data. Furthermore, LieRE is simple to implement and adaptable to modalities, requiring only a tokenizer that also outputs a position in $\mathbb{R}^{d}$ in addition to a standard embedding. In order to aid the reproducibility of our results we will post our code on github.
@@ 1.2.1 Position Encodings
The fact that the original attention mechanism is invariant to the order of tokens has motivated the ongoing development of methods to incorporate positional information into the transformer architecture. We split our literature review into three broad classes of positional encodings: 1) absolute, 2) relative, and 3) contextual.
Absolute encodings generally operate on a per token-level, modifying the embedding of a token to encode the location of the token in the intput. Methods such as sinusoidal and learned absolute encodings add vectors to the input token embedding [vaswani2017attention, devlin2019bert, dosovitskiy2020image]. Absolute position measures position with respect to an absolute reference, such as the start of the text or the top left corner of an image.
Relative position encodings instead encode the relative positions of two tokens. One strategy is to learn an embedding for position deltas which can be incorporated into the attention mechanism [shaw2018self, liu2021swin, liu2022swinv2]. However, this incurs quadratic computational cost in terms of the number of tokens. Rotary Position Encodings (RoPE) avoid this cost by rotating the key and query vectors before the attention inner product. The algebraic properties of the block-diagonal rotation matrices used in RoPE ensures only relative positional information is captured in the attention mechanism [RoFormer2024Su]. RoPE is quite widely used in open source LLMs including the PaLM, Llama and Mixtral models [touvron2023llama, chowdhery2023palm, jiang2024mixtral]. However, RoPE can perform poorly on inference for larger context sizes than the model was trained on. This has spurred an active line of work extending RoPE to longer contexts, work which we review later.
@@ 1.2.2 Extensions of RoPE
One notable one is context extension, which aims to address the fact that RoPE NLP models trained on short documents tend to perform poorly on long documents. Methods like NTK-aware context extension, YaRN and LongRoPE focus on enabling already trained models to handle long context, both with and without finetuning [ding2024longrope, peng2023yarn, tworkowski2024focused, chen2023extending].
Another line of work has been specifically focused on adapting RoPE to image tasks. Both VisionLlama and RoPE-Mixed present relative position encodings inspired by RoPE that are able to encode 2D positional data [chu2024visionllama, heo2024rotary]. The primary difference is that RoPE-Mixed has a learnable component, whereas VisionLlama does not.
@@ 2.1 Lie Groups in the Context of Attention
In this section, we aim to provide a minimal introduction to Lie groups so that the reader is able to understand the mathematical motivations behind LieRE. Lie groups are well studied, especially in the context of representation theory, and standard texts including [fulton2013representation] are able to provide a more extensive introduction to the subject.
In this context, Lie groups are smooth sets of matrices that are closed under matrix multiplication and inversion. For every Lie group, the matrix exponential provides a smooth bijective map from a subset of $\mathbb{R}^{n\times n}$ (hereto referred to as the generator set) to the Lie group. The exponential map is a diffeomorphism and has the following key property for generators $U,V\in\mathbb{R}^{n\times n}$ close together:
$\exp(U-V)=\exp(-V+U)\approx\exp(V)^{-1}\exp(U)$ (1)
@@
$\exp(U-V)=\exp(U)\exp(V)^{-1}\boldsymbol{=}\exp(U)^{-1}\exp(V).$ (2)
Our work examines the tradeoff between using the stronger property in equation (2) or increased capacity and the weaker property (1).
@@ 3.1 Rotations for Relative Positions in Attention
We start by using a learnable linear map to embed the positions as skew-symmetric matrices $P_{i}$ and $P_{j}$ and compute high dimensional rotations as $R_{i}=\exp(P_{i})$ and $R_{j}=\exp(P_{j})$. Learning in the space of skew symmetric matrices allows us to sidestep some of the difficulty that would come from learning on the manifold of rotation matrices.
Recall that, for every pair of tokens, the attention mechanism computes the inner products between their key and query vectors, $k_{i}^{T}q_{j}$, as shown in figure 1. We encode the positions by multiplying by the rotation matrices in the prior part. In particular, we update the keys and queries as $k_{i}^{\prime}=k_{i}R_{i}$ and $q_{j}=q_{j}R_{j}$. This results in an updated inner product of $(R_{i}k_{i})^{T}R_{j}q_{j}=k_{i}R_{i}^{T}R_{j}q_{j}=k_{i}R_{i}^{-1}R_{j}q_{j}.$
Recall that by equation 1, $R_{i}^{-1}R_{j}=\exp(P_{i})^{-1}\exp(P_{j})\approx\exp(P_{j}-P_{i})$. In other words, the inner product automatically computes the relative position encoding. Note that this is the mechanism behind the original RoPE paper. The difference is that LieRE uses the full set of high dimensional rotations and relaxes the constraint that equation 1 holds with equality. We make the description of the method more precise in the next section.
@@ 3.2 LieRE
LieRE modifies the attention mechanism by applying a rotation to the keys and queries before computing their inner products. However, where RoPE applies a block-diagonal rotation matrix with 2D rotation matrices as blocks, LieRE applies a $n\times n$ rotation matrix (Figure 1). Using the fact that every rotation matrix can be represented as the matrix exponential of a skew-symmetric matrix, we parametrize the rotations with generators before the matrix exponential. This skew-symmetric generator is computed via a learned linear map from the positions to the generator space. Let $x\in\mathbb{R}^{n}$ indicate the position of the token, and $A:\mathbb{R}^{d}\rightarrow\text{Skew}_{n}(\mathbb{R})$ be a learnable linear map to the set of skew-symmetric matrices. At the start of every forward pass, we compute the LieRE rotation matrix for each token as
@@
Then, for each attention mechanism, the keys and queries for a particular token are updated as $\mathrm{key^{\prime}}:=R_{LieRE}\cdot\mathrm{key},\mathrm{query^{\prime}}:=R_{LieRE}\cdot\mathrm{query}$ We present attention with LieRE and LieRE-Commute side by side in algorithms 1 and 2.
Algorithm 1 RoPE Attention (prior work for text), rot=rotated
@@
4:         $X_{\text{rot}}^{i}\leftarrow X^{i}\cos p\theta^{i}-X^{i+1}\sin p\theta^{i}$
5:         $X_{\text{rot}}^{i+1}\leftarrow X^{i}\sin p\theta^{i}+X^{i+1}\cos p\theta^{i}$
6:     end for
7:     return $X_{\text{rotated}}$
8:end procedure
9:procedure RoPEAttention($Q,K,V$)
10:     $p\leftarrow\text{tokenPositions}$
@@
12:     $K_{\text{rot}}\leftarrow\textsc{RoPE}(K,p,d)$
13:     $Q_{\text{rot}}\leftarrow\textsc{RoPE}(Q,p,d)$
14:     $\text{Attention}\leftarrow\text{softmax}\left(\frac{Q_{\text{rot}}K_{\text{rot}}^{T}}{\sqrt{d}}\right)V$
15:     return Attention
16:end procedure
Algorithm 2 LieRE Attention (our work), rot=rotated
@@
8:     $K_{\text{rot}}\leftarrow\textsc{LieRE}(K,p,A)$
9:     $Q_{\text{rot}}\leftarrow\textsc{LieRE}(Q,p,A)$
10:     $\text{Attention}\leftarrow\text{softmax}\left(\frac{Q_{\text{rot}}K_{\text{rot}}^{T}}{\sqrt{\text{dim}(K)}}\right)V$
@@
Note that if we apply sparsity structure that enforces 2x2 blocks we recover commutativity. We refer to this configuration and LieRE-Commute. If we then have separate learnable parameters for every layer and head, we recover RoPE-mixed [heo2024rotary].
@@ 4.1 Model and Training Parameters
In order to isolate the effect of changing the position encoding, we use a standard [^1] https://github.com/kentaroy47/vision-transformers-cifar10 transformer backbone modified to be able to switch between relative position encoding types. We use the standard backbone sizes of ViT-Tiny, ViT-B and ViT-L [dosovitskiy2020image]. All experiments use RandAugment [cubuk2020randaugment]. We avoid using pre-trained weights in order to maximize the comparability of results between methods. In order to ensure a fair comparison, we explicitly avoid tuning hyperparameters with LieRE and use the same default hyperparameters for all experiments (Appendix B).
@@ 4.2 Datasets and Tasks
Our experiments are designed to evaluate the efficacy of LieRE as a position encoding across both 2D and 3D data. We evaluate LieRE on the classification of 2D (images) and 3D (videos) data. For 3D data and ImageNet-1k (2D), we focus on accuracy. For CIFAR-100 (2D), where training is less resource intensive, we also evaluate LieRE’s data and training compute efficiency.
@@
For 2D data we evaluate performance on the CIFAR-100 and ImageNet-1k image classification task [krizhevsky2009learning, deng2009imagenet]. We partition our evaluation of performance into four parts. In the first part, we examine accuracy across a range of model architectures on both CIFAR-100 and ImageNet-1k and compare to absolute [dosovitskiy2020image] position encoding and recent related work, RoPE-Mixed [heo2024rotary], VisionLlama embeddings [chu2024visionllama].
In the second part, we take advantage of the relatively modest amount of compute resources necessary to train a model for CIFAR-100 to examine LieREs impact on data efficiency. We also measure training compute efficiency by comparing the number of training steps necessary to achieve a fixed level of validation accuracy.
In the third part, we evaluate the impact of LieRE with various scales. We vary the capacity of the transformer backbone with corresponding to ViT-T, ViT-B and ViT-B, as proposed in [dosovitskiy2020image], the number of learned LieRE generators (one, per-attention-head and per-layer) and the LieRE generator capacity by imposing a block diagonal structure on the generator that allows us to vary the added capacity.
In the fourth part, we measure how much different models depend on the positional information in the image/video by shuffling the patches. A higher accuracy drop with randomly shuffled patches means the model relies more on the positions of the patches during inference.
@@
In this section, we introduce Rotary Position Encodings for 3D data and compare LieRE-based transformers with transformers with RoPE-mixed and absolute encoding similar to the previous section [arnab2021vivit, heo2024rotary]. For the 3D experiments, we examine video classification performance in the UCF101 dataset [soomro2012ucf101]. Again, we did not optimize any hyperparameters for the LieRE model and used the dataloader from [tong2022videomae]. The full set of hyperparameters may be found in appendix B.